# A multi-objective mixed integer linear programming model for supply chain planning of 3D printing

### v2

@@ 4 Problem Definition
Figure 1: Taken from [34], Allocation of parts to 3D printers and sequencing of jobs
Indeed, as demonstrated by Figure 1, we assign the parts to jobs and printers in such a way that customer orders are delivered just on time (minimizing the earliness or tardiness of part delivery), utilizing the printers efficiently given the part orientation optimization and due dates constraints. We call the utilization of a 3D printer efficient if each job utilizes the maximum possible capacity of a 3D printer, i.e., the maximum available surface area in the 3D printer. Furthermore, to simplify the problem, we assume that the required materials for manufacturing parts are of the same type, and the parts are enclosed in cubic shapes to facilitate the calculation of their surface area. The preparation time for each job is combined with the manufacturing time of the parts. The mathematical model can consider one of the three dimensions of the cubic enclosure as the part’s height, depending on how the part is positioned in the 3D printer. Since 3D printers create parts layer by layer, the 3D printer operates at two speeds: volumetric speed and layering speed. Therefore, selecting the appropriate orientation for parts in the 3D printer and consequently choosing the correct part height for timely delivery is highly effective. Because the maximum height in a task determines the layering time, the model must select both an appropriate combination of parts based on their delivery time and an optimal height for them when arranging parts. For example, Figure 2 illustrates a case in which parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers. Figure 3 demonstrates a case in which this problem has been tackled by the correct combination of parts in a job. Note that the printer completes all parts layer by layer.
Figure 2: Taken from [44], combination 1. Parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers
Figure 3: Taken from [44], combination 2. This figure demonstrates a case in which the problem mentioned in Figure 2 has been tackled by the correct combination of parts in a job.
@@
Figure 4: Taken from [44], part orientation
@@ 5 Mathematical Modeling
The developed mathematical model is a multi-objective mixed integer non-linear programming (MOMINLP) model. After the introduction of the sets, parameters, variables, and the MONILP model, we will linearize the model so it can be solved efficiently with known methods.
5.1 Sets
@@ 5.3 Decision Variables
5.4 MOMINLP Model
In the following bi-objective model, we aim to minimize the total tardiness and earliness of parts production and maximize the utilization of machines in each job by the optimal allocation of parts to jobs, jobs to machines, the optimal orientation of parts in each job, and occupying the maximum available area in each machine respecting the constraints.
$z=\min\sum_{i}T_{i}Ct+E_{i}Ce$ (1)
$zz=\min\sum_{j,m}A_{m}y_{jm}-\sum_{i,j,m}A^{{}^{\prime}}_{i}x_{ijm}$ (2)
$\sum_{j,m}x_{ijm}=1$ $\forall i$ (3)
@@ 5.4 MOMINLP Model
The primary objective (1) aims to minimize the costs associated with both earliness and tardiness in part production. Early delivery incurs storage costs, while tardiness can lead to customer dissatisfaction or penalties, with their impacts adjusted by respective cost parameters. The secondary objective function (2) focuses on maximizing machine space utilization by optimizing part allocation to jobs and jobs to machines, considering optimal part orientations and respecting all constraints. Constraints (3) ensure each part is assigned to exactly one job on one machine. Part assignment to a job on a machine is contingent upon the job being allocated to that machine, governed by constraints (5). Constraints (5) to (9) dictate part orientation and placement within the printer, ensuring only one base side of the part is chosen, determining its vertical dimension, calculating base area, limiting part height within printer constraints, and enforcing printer capacity limits. Constraint (10) selects the maximum height among parts within a job, while (11) computes job processing times on machines without considering setup times. Printing time vertically and volumetrically is computed accordingly. Constraints (12) ensure job completion times on a machine follow proper sequencing, with subsequent job completion times greater than or equal to previous ones plus respective processing times. Initial job completion times on each machine are established by (13). Constraint (14) links part completion times to their respective job completion times. Tardiness (15) and earliness (16) for each part are computed, ensuring non-negative earliness. Finally, (17) enforces proper job sequencing on the same machine.
@@ 5.5 Linearization
Additionally, the following set of constraints will ensure that $lp_{ijm}=A^{{}^{\prime}}_{i}$ if $x_{ijm}=1$ and 00 otherwise in a linear fashion:
@@
Additionally, the following set of constraints will ensure that $lp^{{}^{\prime}}_{ijm}=C_{jm}$ if $x_{ijm}=1$ and 00 otherwise in a linear fashion:
@@ 6.2 Scenario Analysis
666688881010101012121212141414141616161618181818202020200020202020404040406060606080808080Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
@@
6666888810101010121212121414141416161616181818182020202000101010102020202030303030Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
@@
6666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202000202020204040404060606060Number of partsSum of earliness and tardinessScenario 1 (Single Machine)Scenario 1 (Two Machines)Scenario 2 (Single Machine)Scenario 2 (Two Machines)
@@
According to Figure 8, the total delay and lead time for parts in both scenarios, with two machines, have decreased by at least $50$ compared to having only one machine. This reduction is greater for the first scenario than for the second scenario. Specifically, the decrease in the scenario is over $70$ for the first $17$ parts, whereas in the second scenario, this reduction only occurs for the first $10$ parts. An important point to note is that with an increase in the number of parts, the reduction percentage decreases to $50$ for both scenarios. This implies that increasing the number of parts may lead the first scenario also to face machine space constraints, making some desired heights unattainable due to space limitations. However, it is undeniable that the first scenario has performed significantly better than the second scenario. If the number of parts is very high, the model may inevitably choose undesirable heights for parts due to machine space constraints. Nonetheless, the second scenario will not perform better than the first scenario as the second scenario is a feasible solution for the model and sub-optimal generally.
@@ 6.3 Sensitivity Analysis
Table 4: Dimensions of parts for sensitivity analysis of layering time parameter ($Vt$)
@@
$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$00200200200200400400400400Layer production time (h/mm)Sum of earliness and tardiness (h)Scenario 1Scenario 2
@@
222244446666888810101010121212121414141400100100100100200200200200300300300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
@@
222244446666888810101010121212121414141400100100100100200200200200300300300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
@@
22224444666688881010101012121212141414140020202020404040406060606080808080100100100100Number of partsSum of earliness and tardiness (h)Scenario 1, $A_{m}=4000$ $mm^{2}$cenario 1, $A_{m}=9000$ $mm^{2}$
@@ 7 Conclusions
3D printing stands at the forefront of addressing global challenges posed by population growth and resource demands by offering efficient, sustainable production systems. Unlike traditional methods, 3D printing excels in minimizing material usage and environmental impact while enabling customized, on-demand production across industrial, medical, and aerospace sectors. Efficient production planning and scheduling are pivotal in maximizing the advantages of 3D printing. Integrating just-in-time principles with the optimization of part orientation and machine utilization presents a progressive approach to enhancing operational efficiency and cost-effectiveness. Our study contributes to this area by introducing a multi-objective mixed integer linear programming (MOMILP) model designed to optimize 3D printing production schedules, minimizing earliness and tardiness and maximizing 3D printer utilization. This project addresses the complex task of allocating parts with diverse characteristics to jobs and scheduling these jobs across multiple machines. With a focus on minimizing delay and lead time while maximizing machine utilization, the study explores three distinct part placement scenarios. The mathematical model, designed with dual objectives, aims to achieve Pareto optimal solutions due to the inherent trade-offs involved. An innovative aspect of the model involves investigating how selecting a part’s height dimension affects its delivery time, simplifying the analysis to emphasize time-related efficiencies in 3D printing operations. This paper focuses on the allocation of parts with various characteristics to jobs and scheduling these jobs, considering the orientation optimization of parts in machines. Given the dual objectives of the mathematical model presented, a universally optimal solution is not expected. Therefore, a numerical example was provided and solved using the epsilon-constraint method, resulting in the Pareto optimal curve composed of several achievable solutions. One of the ideas proposed in the model is to examine the effect of selecting one dimension of the part as its height on the delivery time. Key findings from our analysis are as follows:
@@
Future endeavors could incorporate parts made from various materials, requiring strategic production planning across different machines. Exploring printers with multiple nozzles operating at varying speeds presents new scheduling challenges and opportunities for efficiency gains. Additionally, leveraging sophisticated heuristic methods could optimize solutions for large-sized problems, acknowledging the problem’s NP-hard complexity. Also, integrating cost-related metrics into objective functions would provide insights into overall expense management and sustainability. This project represents a foundational step towards advancing production scheduling in 3D printing, particularly focusing on time efficiency as a pivotal factor for industrial adoption and scalability.

### v4

@@ 4 Problem Definition
Figure 1: Adapted from [34], allocation of parts to 3D printers and sequencing of jobs
Indeed, as demonstrated by Figure 1, we assign the parts to jobs and printers in such a way that customer orders are delivered just on time (minimizing the earliness or tardiness of part delivery), utilizing the printers efficiently given the part orientation optimization and due dates constraints.
We call the utilization of a 3D printer efficient if each job utilizes the maximum possible capacity of a 3D printer, i.e., the maximum available surface area in the 3D printer. Furthermore, to simplify the problem, we assume that the required materials for manufacturing parts are of the same type, and the parts are enclosed in cubic shapes to facilitate the calculation of their surface area. The preparation time for each job is combined with the manufacturing time of the parts.
The mathematical model can consider one of the three dimensions of the cubic enclosure as the part’s height, depending on how the part is positioned in the 3D printer. Since 3D printers create parts layer by layer, the 3D printer operates at two speeds: volumetric speed and layering speed. Therefore, selecting the appropriate orientation for parts in the 3D printer and consequently choosing the correct part height for timely delivery is highly effective.
Because the maximum height in a task determines the layering time, the model must select both an appropriate combination of parts based on their delivery time and an optimal height for them when arranging parts. For example, Figure 2 illustrates a case in which parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers. Figure 3 demonstrates a case in which this problem has been tackled by the correct combination of parts in a job. Note that the printer completes all parts layer by layer.
Figure 2: Adapted from [44], combination 1. Parts $P_{2}$ and $P_{4}$ determine the completion of the job while the other parts have been completed earlier. This can cause delays and inefficiency in the utilization of the printers.
Figure 3: Adapted from [44], combination 2. This figure demonstrates a case in which the problem mentioned in Figure 2 has been tackled by the correct combination of parts in a job.
@@
Figure 4: Adapted from [44], part orientation
@@ 5 Mathematical Modeling
After the introduction of the sets, parameters, variables, and the MONILP model, we will linearize the model so it can be solved efficiently with known methods.
Sets
@@
MOMINLP Model
In the following bi-objective model, we aim to minimize the total tardiness and earliness of parts production and maximize the utilization of machines in each job by the optimal allocation of parts to jobs, jobs to machines, the optimal orientation of parts in each job, and occupying the maximum available area in each machine, respecting the constraints.
$z=\min_{T,E}$ $\sum_{i}T_{i}\cdot Ct+E_{i}\cdot Ce$ (1)
$zz=\min_{y,x}$ $\sum_{j,m}A_{m}y_{jm}-\sum_{i,j,m}A^{{}^{\prime}}_{i}x_{ijm}$ (2)
$s.t.$ $\sum_{j,m}x_{ijm}=1$ $\forall i$ (3)
@@
The primary objective (1) aims to minimize the costs associated with both earliness and tardiness in part production. Early delivery incurs storage costs, while tardiness can lead to customer dissatisfaction or penalties, with their impacts adjusted by respective cost parameters.
The secondary objective function (2) focuses on maximizing machine space utilization by optimizing part allocation to jobs and jobs to machines, considering optimal part orientations and respecting all constraints.
Constraints (3) ensure each part is assigned to exactly one job on one machine. Part assignment to a job on a machine is contingent upon the job being allocated to that machine, governed by constraints (5).
Constraints (5) to (9) dictate part orientation and placement within the printer, ensuring only one base side of the part is chosen, determining its vertical dimension, calculating base area, limiting part height within printer constraints, and enforcing printer capacity limits.
Constraint (10) selects the maximum height among parts within a job, while (11) computes job processing times on machines without considering setup times. Printing time vertically and volumetrically is computed accordingly.
Constraints (12) ensure job completion times on a machine follow proper sequencing, with subsequent job completion times greater than or equal to previous ones plus respective processing times. Initial job completion times on each machine are established by (13). Constraint (14) links part completion times to their respective job completion times. Tardiness (15) and earliness (16) for each part are computed, ensuring non-negative earliness. Finally, (17) enforces proper job sequencing on the same machine.
@@ 5.1 Linearization
Additionally, the following set of constraints will ensure that $lp_{ijm}=A^{{}^{\prime}}_{i}$ if $x_{ijm}=1$ and 0 otherwise in a linear fashion:
@@
Additionally, the following set of constraints will ensure that $lp^{{}^{\prime}}_{ijm}=C_{jm}$ if $x_{ijm}=1$ and 0 otherwise in a linear fashion:
@@ 6.2 Scenario Analysis
668810101212141416161818202002020404060608080Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
@@
66881010121214141616181820200101020203030Number of partsSum of earliness and tardinessFirst ScenarioSecond Scenario
@@
66881010121214141616181820200202040406060Number of partsSum of earliness and tardinessScenario 1 (Single Machine)Scenario 1 (Two Machines)Scenario 2 (Single Machine)Scenario 2 (Two Machines)
@@
According to Figure 8, the total delay and lead time for parts in both scenarios, with two machines, have decreased by at least $50$ % compared to having only one machine. This reduction is greater for the first scenario than for the second scenario. Specifically, the decrease in the scenario is over $70$ % for the first $17$ parts, whereas in the second scenario, this reduction only occurs for the first $10$ parts. An important point to note is that with an increase in the number of parts, the reduction percentage decreases to $50$ % for both scenarios. This implies that increasing the number of parts may lead the first scenario also to face machine space constraints, making some desired heights unattainable due to space limitations. However, it is undeniable that the first scenario has performed significantly better than the second scenario. If the number of parts is very high, the model may inevitably choose undesirable heights for parts due to machine space constraints. Nonetheless, the second scenario will not perform better than the first scenario as the second scenario is a feasible solution for the model and sub-optimal generally.
@@ 6.3 Sensitivity Analysis
Table 4: Dimensions of parts for sensitivity analysis of layering time parameter (VtVt)
@@
$10^{-6}$$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$0200200400400Layer production time (h/$m$)Sum of earliness and tardiness (h)Scenario 1Scenario 2
@@
224466881010121214140100100200200300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
@@
224466881010121214140100100200200300300Number of partsSum of earliness and tardiness (h)Scenario 1Scenario 2
@@
2244668810101212141402020404060608080100100Number of partsSum of earliness and tardiness (h)Scenario 1, $A_{m}=4000$ $mm^{2}$Scenario 1, $A_{m}=9000$ $mm^{2}$
@@ 7 Conclusions
3D printing stands at the forefront of addressing global challenges posed by population growth and resource demands by offering efficient, sustainable production systems. Unlike traditional methods, 3D printing excels in minimizing material usage and environmental impact while enabling customized, on-demand production across industrial, medical, and aerospace sectors. Efficient production planning and scheduling are pivotal in maximizing the advantages of 3D printing. Integrating just-in-time principles with the optimization of part orientation and machine utilization presents a progressive approach to enhancing operational efficiency and cost-effectiveness.
Our study contributes to this area by introducing a multi-objective mixed integer linear programming model designed to optimize 3D printing production schedules, minimizing earliness and tardiness and maximizing 3D printer utilization.
This project addresses the complex task of allocating parts with diverse characteristics to jobs and scheduling these jobs across multiple machines. With a focus on minimizing delay and lead time while maximizing machine utilization, the study explores three distinct part placement scenarios. The mathematical model, designed with dual objectives, aims to achieve Pareto optimal solutions due to the inherent trade-offs involved. An innovative aspect of the model involves investigating how selecting a part’s height dimension affects its delivery time, simplifying the analysis to emphasize time-related efficiencies in 3D printing operations. This paper focuses on the allocation of parts with various characteristics to jobs and scheduling these jobs, considering the orientation optimization of parts in machines. Given the dual objectives of the mathematical model presented, a universally optimal solution is not expected. Therefore, a numerical example was provided and solved using the epsilon-constraint method, resulting in the Pareto optimal curve composed of several achievable solutions.
One of the ideas proposed in the model is to examine the effect of selecting one dimension of the part as its height on the delivery time. Key findings from our analysis are as follows: